# All for One, and One for All: UrbanSyn Dataset, the third Musketeer of Synthetic Driving Scenes

### v1

@@ III-E UrbanSyn brings diversity
GTAV was acquired from a video game, bringing all the well-known limitations related to real-time simulation (e.g., limited degrees of graphics realism, and a limited number of object instances present at once in the scene). Additionally, GTAV suffers a clear bias towards American-style street maps and cars. Synscapes was designed to closely resemble the Cityscapes dataset in structure and content as stated by their authors [2]. Unlike GTAV, the scenarios generated in Synscapes are biased toward German cities. It is based on small maps, making use of a limited variety of materials, simplified procedural roads and buildings, missing real-world imperfections, and thus striking CGI uncanny perfection. Moreover, images tend to be overpopulated for all classes. On the other hand, Synscapes offered higher image quality by utilizing unbiased path-tracing rendering.
UrbanSyn has been conceived for diversity since its conception, not being tailored for any specific real-world dataset. Having in mind that more synthetic data does not necessarily guarantee a better model performance [67], we defined a set of clear requirements for the newly proposed dataset: i) it should achieve high graphics realism, at the level of Synscapes or beyond, to help to reduce synth-to-real visual domain shift; ii) the images should depict scenarios unseen in previous datasets, including real-world imperfections; iii) we aimed for semi-procedurally generated content seamlessly integrated with high-quality assets, so to get the better of both worlds in terms of realism and the covered variety. The achievement of these requirements has made UrbanSyn a great complement for GTAV and Synscapes. Thus, it makes sense to use UrbanSyn alone or as part of the Three Musketeers.
@@ IV Experimental results
In this paper, we highlight the usefulness of UrbanSyn by focusing on one of its ground truths, namely the pixel-level class labels. These labels are crucial for conducting synth-to-real UDA to develop semantic segmentation models without human labeling. We conduct comprehensive experiments in this section to demonstrate the importance of these labels. We begin by presenting the datasets and evaluation metrics used in our experiments. Next, we provide additional implementation details. Finally, we present our experiments and a discussion of the results obtained.
IV-A Datasets and evaluation
@@
The source (synthetic) datasets in common use to address synth-to-real UDA for semantic segmentation are GTAV [1], Synscapes [2], and SYNTHIA [4]. GTAV and Synscapes have semantic labeling for the 19 classes of Cityscapes [18], while this is not the case for SYNTHIA. Thus, since we are especially interested in a multi-source setting, in this paper, we define the Three Musketeers dataset as consisting of GTAV, Synscapes, and UrbanSyn. GTAV is composed of 24,966 images with a resolution of pixels directly obtained from the render engine of the videogame GTA V. Synscapes is composed of 25,000 images with a resolution of pixels of urban scenes, obtained by using a physic-based rendering pipeline. UrbanSyn is composed of 7,539 images with a resolution of pixels.
@@
As target (real-world) datasets, we not only consider Cityscapes [18] as in most previous literature but also the more challenging BDD100K [19] and Mapillary Vistas [20] datasets. Cityscapes is well-known for its semantic segmentation challenge, composed of 2,975 images for training and 500 for validation. The images have a resolution of pixels and were obtained from different cities in Germany. BDD100K contains images taken from different vehicles in different USA cities. It is divided into 7,000 images for training and 1,000 for validation, with a resolution of pixels. As in [22], for training purposes, we only use daytime images without heavy occlusions and favorable weather, resulting in a subset of 1,777 images. We consider the validation set as it is. Mapillary Vistas is composed of a large number of images from around the world. These images were obtained with different camera devices generating a large variety of resolutions and aspect ratios. Some images were captured onboard and others from personal devices at street level. As in [22], we only consider images with an aspect ratio of 4:3, which comprises more than 75% of all the dataset images. This results in 14,716 images for training and 1,617 for validation.
@@
For each model that we train, we assess its accuracy in the validation set of each real-world dataset according to the 19 official classes defined for Cityscapes. As is common practice, other classes are ignored. Model training in the synth-to-real UDA setting is performed by considering the synthetic datasets and the training set of each real-world dataset. The semantic labels of these real-world training images are ignored. These labels are just considered to train reference models, i.e., those trained and validated on the same real-world dataset domain. As is standard, to assess segmentation accuracy, we use the intersection-over-union metric [68], i.e., , where TP, FP, and FN refer to true positives, false positives, and false negatives, respectively. IoU can be computed per class while using a mean IoU (mIoU) to consider all the classes at once. We repeat each experiment three times and report its average and standard deviation.
@@
| DeepLabV3+ | Cityscapes (reference) | std |
@@
| DeeplabV3+ | Mapillary Vistas (reference) | std |
IV-B Implementation details
We perform synth-to-real LAB space alignment as a best practice preprocessing which significantly reduces visual domain gaps [54, 22]. Here, combining different datasets means aggregating them to become a single but multi-source synthetic domain. During the selection of images to compose training mini-batches, we take into account that different sources have different numbers of images to ensure the same probability of selection per image, regardless of the image source.
@@ IV-B Implementation details
We consider models trained on synthetic data as baselines. All the baselines involving DeepLabV3+ are trained for 90K iterations with SGD optimizer with a starting learning rate of 0.002, and momentum of 0.9. For Cityscapes, BDD100K, and Mapillary Vistas, we crop the training images to pixels, , and ; and batch sizes are set to 8, 16, and 4 images, respectively. SegFormer baselines are trained for 160K iterations with a batch size of 2 in all configurations, using the same hyper-parameters, optimizer, and scheduler reported in [17], with the same cropping adjustments as for the DeepLabV3+ baselines. For co-training, we use the hyper-parameters reported in [22], except for the batch size that we have been able to increase to the same values used here for the DeepLabV3+ baselines. The crop sizes used in [22] and here are the same. For HRDA, we use the settings stated in [21], except for crop sizes which are set to match our SegFormer baselines. In addition, since our focus is on using multiple synthetic sources, following the instructions of the HRDA framework, we have deactivated the thing-class ImageNet feature distance (FD) inherent to DAFormer-based approaches, which is tailored for the GTAV-to-Cityspaces UDA case.
IV-C Results analysis
@@
This analysis relies on the results reported in Tables I, II, and III, which correspond to Cityscapes, BDD100K, and Mapillary Vistas, respectively. Each table has a block for approaches based on DeepLabV3+ and another for those based on SegFormer, each including sub-blocks (separated by horizontal lines) for baselines, UDA, and the block reference (training and testing in the same real-world domain).
@@
Comparing the individual performance of UrbanSyn, GTAV, and Synscapes, we see that the best per-class accuracy is not always given by the same dataset. However, in terms of mIoU, UrbanSyn is the best performing in all cases (always for DeepLabV3+; with Cityscapes and Mapillary Vistas for SegFormer) but one (with BDD100K for SegFormer, on pair with GTAV). The same outcome is observed when applying single-source synth-to-real UDA. In terms of baselines, we see that by adding UrbanSyn to GTAV and Synscapes, the mIoU increases significantly: points for SegFormer/Cityscapes, for DeepLabV3+/Cityscapes, for SegFormer/BDD100K, for DeepLabV3+/BDD100K, for SegFormer/Mapillary Vistas, for DeepLabV3+/Mapillary Vistas. Finally, we see that Musketeers is also beneficial for synth-to-real UDA. For Cityscapes, HRDA reaches mIoU=, which is just points behind its reference; while co-training reaches mIoU=, just points behind its reference. For BDD100K, HRDA reaches mIoU=, points behind its reference; while co-training reaches mIoU=, points behind its reference. For Mapillary Vistas, HRDA reaches mIoU=, points behind its reference; while co-training reaches mIoU=, points behind its reference.
@@
It is noticeable that Musketeers with HRDA (without ImageNet distance loss) reach mIoU= for Cityscapes, which is the new synth-to-real state-of-the-art to the best of our knowledge. The mIoU= reached by co-training is points better than the results reported in [22]. For the case of BDD100K and Mapillary Vistas, co-training improves and points the results reported in [22].
@@
Figure 4 illustrate how UrbanSyn is helping to provide stronger baselines as part of the Musketeers, no matter if we use DeepLabV3+ or SegFormer. For Cityscapes, UrbanSyn helps to better separate the road from the sidewalk; for BDD100K, it helps to segment better the vehicles; and, for Mapillary Vistas, traffic signs are better-segmented thanks to UrbanSyn. In addition, looking at the qualitative results of the different validation sets, what we observe is that by using UrbanSyn, we improve the segmentation of objects that occupy a small image area (e.g., traffic signs and lights, poles, pedestrians, etc), which is usually the case when these objects are at mid and large distances from the ego-vehicle.
@@
IV-D Erroneous human labels
During the qualitative analysis of the results, we observed labeling errors in the ground truth of the real-world datasets. For instance, the BDD100K validation set only contains seven images (out of 1,000 images) with labels of the Train class. In five of these, the samples are extremely challenging (Figure 5), while in the other two, the labels are just wrong (Figure 6). Analyzing the BDD100K training set, we find only 27 images with labels of the Train class (out of 7,000 images), where the labels are wrong in 13 of them. All this explains the poor results seen in Table II for the Train class. As can be in Figures 7 and 8, Cityscapes also have some errors and inconsistencies regarding human labeling. In Figure 7, we see how the co-training method properly labels a traffic light and a traffic sign, while humans did not label them. Since this image is from the validation set of Cityscapes, the self-labels are considered wrong. In the same image, a rider on a bicycle is labeled as a fully visible Person by humans, while the upper part of the person is labeled as a Rider by co-training, and the rest is labeled as a Bicycle. Again, we think that the self-labels represent better the semantics of this image patch. In Figure 8, we can see how different criteria have been used to label Fence instances in different images. 9 shows another example of inconsistency, in this case from a Mapillary Vistas image. We can see how the closed Traffic Light instance has been fully labeled as such (lightbox and its support), while in the further away ones, only the light boxes are labeled as Traffic Light.

### v2

@@ III-E UrbanSyn brings diversity
GTAV was acquired from a video game, bringing all the well-known limitations related to real-time simulation (e.g., limited degrees of graphics realism, and a limited number of object instances present at once in the scene). Additionally, GTAV suffers a clear bias towards American-style street maps and cars. Synscapes was designed to closely resemble the Cityscapes dataset in structure and content as stated by their authors [2]. Unlike GTAV, the scenarios generated in Synscapes are biased toward German cities. It is based on small maps, making use of a limited variety of materials, simplified procedural roads and buildings, missing real-world imperfections, and thus striking CGI uncanny perfection. Moreover, images tend to be overpopulated for all classes lacking in realism on scene composition. On the other hand, Synscapes offered higher image quality by utilizing unbiased path-tracing rendering.
UrbanSyn has been conceived for diversity since its conception, not being tailored for any specific real-world dataset. Having in mind that more synthetic data does not necessarily guarantee a better model performance [72], we defined a set of clear requirements for the newly proposed dataset: i) it should achieve high graphics realism, at the level of Synscapes or beyond, to help to reduce synth-to-real visual domain shift; ii) the images should depict scenarios unseen in previous datasets, including real-world imperfections; iii) we aimed for semi-procedurally generated content seamlessly integrated with high-quality assets, so to get the better of both worlds in terms of realism and the covered variety. The achievement of these requirements has made UrbanSyn a great complement for GTAV and Synscapes. Thus, it makes sense to use UrbanSyn alone or as part of the Three Musketeers. Figure 4 details per-class statistics of pixel occupancy and percentage of images of the Three Musketeers. UrbanSyn is well-balanced and a trade-off between GTAV and Synscapes. On the one hand, it provides a more balanced variability and quantity than GTAV in classes such as Rider, Bus, Train, Motorcycle, and Bicycle. On the other hand, UrbanSyn maintains the scene realism and coherence, avoiding Synscapes practices to overpopulate all images always with every class on them (e.g., Synscapes have around 70% of images that contain class Train, unrealistic in a real-world urban scenario.)
@@ IV Experimental results
In this paper, we highlight the usefulness of UrbanSyn by focusing on one of its ground truths, namely the pixel-level class labels. These labels are crucial for conducting synth-to-real UDA to develop semantic segmentation models without human labeling. We conduct comprehensive experiments in this section to demonstrate the importance of these labels.In the rest of this section, we begin by explaining the standard experimental methodology applied to evaluate synth-to-real UDA settings, for either assessing new UDA protocols or new synthetic datasets as is our purpose in this paper. Next, we detail the datasets and evaluation metrics used in our experiments. After, we provide additional implementation details. Finally, we present the experiments and discuss the results obtained.
IV-A Experimental methodology
Figure 5 illustrates the standard experimental evaluation methodology used to assess synth-to-real UDA procedures, which we also adhere to in this paper. It is assumed that we have a deep model to be adapted, a source domain in the form of automatically labeled synthetic data, , and a target domain in the form of real-world unlabeled training data, . In practice, to perform experiments, the real-world domain is composed of a training and a testing set, and respectively. These are used by the research community for different types of tasks. For instance, in the case of synth-to-real UDA, the set plays the role of by ignoring its (human-provided) labels. Then, a standard training of the deep model using is used as upper-bound reference for the synth-to-real UDA under evaluation; while an analogous training based on is used as lower-bound reference. The goal of a synth-to-real UDA method is to use and to obtain a similar (or better) performing deep model than if is used as the training set. This process involves self-labeling . To assess model accuracy irrespective of its training, and a standard evaluation procedure are used. The assumption is that good synth-to-real UDA procedures will be able to address the training of deep models for new unlabeled real-world datasets at least using the same synthetic data or improved versions of it.
This general methodology translates into specific experiments to assess the goodness of UrbanSyn. Table III summarizes them. We run this table of experiments twice, i.e., to cover two significantly different deep models for semantic segmentation and their respective UDA procedures, namely, DeepLabV3+ (pure CNN-based model) and SegFormer MiT-B5 (Transformer-based model) with Co-training and HRDA respective UDA procedures. Comparing experiments with corresponding , where and , we can see domain gaps in accuracy for the supervised settings. Note that in the experiments, real-world human-labeled images are used for training and there is no domain gap, thus, the resulting accuracy can be taken as an upper-bound reference. In the experiments, synthetic-only automatically labeled images are used for training and there is a domain gap with the real-world images used for testing, thus, the resulting accuracy can be taken as a lower-bound reference or baseline. We can see how large the domain gap between given and is by comparing the accuracy resulting from the and experiments. In the UDA setting, we can see the contribution of the synthetic data by comparing the accuracy of experiments to and , varying and the UDA method, where the aim is to approach (or surpass) the accuracy of after significantly improving . If we restrict to a single synthetic dataset (here ), we can compare their individual contributions. If comes from the combination of different synthetic datasets, we can evaluate if they are complementary to the UDA task.
According to this standard setting of experiments, we will see in section IV-D that UrbanSyn allows to significantly increase the accuracy of different state-of-the-art synth-to-real UDA procedures for semantic segmentation (DeepLabV3+/Co-training, SegFormer MiT-B5/HRDA), in single-source regime–individually comparing UrbanSyn (U) to GTAV (G) and Synscapes (S)–and in a multi-source regime–comparing the use of GTAV+Synscapes+UrbanSyn (GSU) to only using GTAV+Synscapes (GS)–, and working with different real-world target domains–Cityscapes (C), BDD10K (B), Mapillary Vistas (M)–.
Experiment | from … | from … | from … | Where? | |
Type ID | |||||
Real-to-real supervised training (same domain) | |||||
C-C-Sup | - | Cityscapes | - | Cityscapes | Table IV |
B-B-Sup | BDD100K | BDD100K | Table V | ||
M-M-Sup | Mapillary | Mapillary | Table VI | ||
Synth-to-real supervised training (domain gap, no UDA) | |||||
G-C-Sup | GTAV | - | - | Cityscapes | Table IV |
G-B-Sup | BDD100K | Table V | |||
G-M-Sup | Mapillary | Table VI | |||
S-C-Sup | Synscapes | - | - | Cityscapes | Table IV |
S-B-Sup | BDD100K | Table V | |||
S-M-Sup | Mapillary | Table VI | |||
U-C-Sup | UrbanSyn | - | - | Cityscapes | Table IV |
U-B-Sup | BDD100K | Table V | |||
U-M-Sup | Mapillary | Table VI | |||
GS-C-Sup | GTAV | - | - | Cityscapes | Table IV |
GS-B-Sup | + | BDD100K | Table V | ||
GS-M-Sup | Synscapes | Mapillary | Table VI | ||
GSU-C-Sup | GTAV | - | - | Cityscapes | Table IV |
GSU-B-Sup | + Synscapes | BDD100K | Table V | ||
GSU-M-Sup | + Urbansyn | Mapillary | Table VI | ||
Synth-to-real UDA training (domain gap, self-labeling based UDA) | |||||
G-C-UDA | GTAV | - | Cityscapes | Cityscapes | Table IV |
G-B-UDA | BDD100K | BDD100K | Table V | ||
G-M-UDA | Mapillary | Mapillary | Table VI | ||
S-C-UDA | Synscapes | - | Cityscapes | Cityscapes | Table IV |
S-B-UDA | BDD100K | BDD100K | Table V | ||
S-M-UDA | Mapillary | Mapillary | Table VI | ||
U-C-UDA | UrbanSyn | - | Cityscapes | Cityscapes | Table IV |
U-B-UDA | BDD100K | BDD100K | Table V | ||
U-M-UDA | Mapillary | Mapillary | Table VI | ||
GSU-C-UDA | GTAV | - | Cityscapes | Cityscapes | Table IV |
GSU-B-UDA | + Synscapes | BDD100K | BDD100K | Table V | |
GSU-M-UDA | + Urbansyn | Mapillary | Mapillary | Table VI | |
IV-B Datasets and evaluation
@@
The source (synthetic) datasets in common use to address synth-to-real UDA for semantic segmentation are GTAV [1], Synscapes [2], and SYNTHIA [4]. GTAV and Synscapes have semantic labeling for the 19 classes of Cityscapes [25], while this is not the case for SYNTHIA. Thus, since we are especially interested in a multi-source setting, in this paper, we define the Three Musketeers dataset as consisting of GTAV, Synscapes, and UrbanSyn. GTAV is composed of 24,966 images with a resolution of pixels directly obtained from the render engine of the videogame GTA V. Synscapes is composed of 25,000 images with a resolution of pixels of urban scenes, obtained by using a physics-based rendering pipeline. UrbanSyn is composed of 7,539 images with a resolution of pixels.
@@
As target (real-world) datasets, we not only consider Cityscapes [25] as in most previous literature but also the more challenging BDD100K [26] and Mapillary Vistas [27] datasets. Cityscapes is well-known for its semantic segmentation challenge, comprising 2,975 images for training and 500 for validation. The images have a resolution of pixels and were obtained from different cities in Germany. BDD100K contains images taken from different vehicles in different USA cities. It is divided into 7,000 images for training and 1,000 for validation, with a resolution of pixels. As in [16], for training purposes, we only use daytime images without heavy occlusions and favorable weather, resulting in a subset of 1,777 images. We consider the validation set as it comes. Mapillary Vistas is composed of a large number of images from around the world. These images were obtained with different camera devices, generating a large variety of resolutions and aspect ratios. Some images were captured onboard, and others were taken from personal devices at the street level. As in [16], we only consider images with an aspect ratio of 4:3, which comprises more than 75% of all the dataset images. This results in 14,716 images for training and 1,617 for validation.
@@
For each model that we train, we assess its accuracy in the validation set of each real-world dataset according to the 19 official classes defined for Cityscapes. As is common practice, other classes are ignored. Model training in the synth-to-real UDA setting is performed by considering the synthetic datasets and the training set of each real-world dataset. The semantic labels of these real-world training images are ignored. These labels are just considered to train reference models, i.e., those trained and validated on the same real-world dataset domain. As is standard, to assess segmentation accuracy, we use the intersection-over-union metric [73], i.e., , where TP, FP, and FN refer to true positives, false positives, and false negatives, respectively. Given a pixel with ground truth class and predicted class , if it counts as TP for class , if it counts as FN for class and as FP for class . Since IoU is computed per class, their mean (mIoU) is used to consider all the classes at once. We repeat each experiment three times and report its average and standard deviation.
@@
DeepLabV3+ | Cityscapes (reference) | std | ||||||||||||||||||||
@@
DeeplabV3+ | Mapillary Vistas (reference) | std | ||||||||||||||||||||
IV-C Implementation details
We perform synth-to-real LAB space alignment as a best practice preprocessing that significantly reduces visual domain gaps [15, 16]. Here, combining different datasets means aggregating them to become a single but multi-source synthetic domain. During the selection of images to compose training mini-batches, we take into account that different sources have different numbers of images to ensure the same probability of selection per image, regardless of the image source.
@@ IV-C Implementation details
We consider models trained on synthetic data as baselines. All the baselines involving DeepLabV3+ are trained for 90K iterations with SGD optimizer with a starting learning rate of 0.002 and momentum of 0.9. For Cityscapes, BDD100K, and Mapillary Vistas, we crop the training images to pixels, , and ; and batch sizes are set to 8, 16, and 4 images, respectively. SegFormer baselines are trained for 160K iterations with a batch size of 2 in all configurations, using the same hyper-parameters, optimizer, and scheduler reported in [24], with the same cropping adjustments as for the DeepLabV3+ baselines. For co-training, we use the hyper-parameters reported in [16], except for the batch size that we have been able to increase to the same values used here for the DeepLabV3+ baselines. The crop sizes used in [16] and here are the same. For HRDA, we use the settings stated in [28], except for crop sizes, which are set to match our SegFormer baselines. In addition, since our focus is on using multiple synthetic sources, following the instructions of the HRDA framework, we have deactivated the thing-class ImageNet feature distance (FD) inherent to DAFormer-based approaches, which is tailored for the GTAV-to-Cityspaces UDA case.
IV-D Results analysis
@@
This analysis relies on the results reported in Tables IV, V, and VI, which correspond to Cityscapes, BDD100K, and Mapillary Vistas, respectively. Each table has a block for approaches based on DeepLabV3+ and another for those based on SegFormer, each including sub-blocks (separated by horizontal lines) for baselines, UDA, and the block reference (training and testing in the same real-world domain).
@@
Comparing the individual performance of UrbanSyn, GTAV, and Synscapes, we see that the best per-class accuracy is not always given by the same dataset. However, in terms of mIoU, UrbanSyn is the best performing in all cases (always for DeepLabV3+; with Cityscapes and Mapillary Vistas for SegFormer) but one (with BDD100K for SegFormer, in pair with GTAV). The same outcome is observed when applying single-source synth-to-real UDA. In terms of baselines, we see that by adding UrbanSyn to GTAV and Synscapes, the mIoU increases significantly: points for SegFormer/Cityscapes, for DeepLabV3+/Cityscapes, for SegFormer/BDD100K, for DeepLabV3+/BDD100K, for SegFormer/Mapillary Vistas, for DeepLabV3+/Mapillary Vistas. Finally, we see that Musketeers is also beneficial for synth-to-real UDA. For Cityscapes, HRDA reaches mIoU=, which is just points behind its reference, while co-training reaches mIoU=, just points behind its reference. For BDD100K, HRDA reaches mIoU=, points behind its reference, while co-training reaches mIoU=, points behind its reference. For Mapillary Vistas, HRDA reaches mIoU=, points behind its reference, while co-training reaches mIoU=, points behind its reference.
@@
It is noticeable that Musketeers with HRDA (without ImageNet distance loss) reach mIoU= for Cityscapes, which is the new synth-to-real state-of-the-art to the best of our knowledge. The mIoU= reached by co-training is points better than the results reported in [16]. For the case of BDD100K and Mapillary Vistas, co-training improves and points from the results reported in [16].
@@
Figure 6 illustrates how UrbanSyn is helping to provide stronger baselines as part of the Musketeers, no matter if we use DeepLabV3+ or SegFormer. For Cityscapes, UrbanSyn helps to better separate the road from the sidewalk; for BDD100K, it helps to segment the vehicles better; and, for Mapillary Vistas, traffic signs are better-segmented thanks to UrbanSyn. In addition, looking at the qualitative results of the different validation sets, what we observe is that by using UrbanSyn, we improve the segmentation of objects that occupy a small image area (e.g., traffic signs and lights, poles, pedestrians, etc), which is usually the case when these objects are at mid and large distances from the ego-vehicle.
@@
IV-E Erroneous human labels
During the qualitative analysis of the results, we observed labeling errors in the ground truth of the real-world datasets. For instance, the BDD100K validation set only contains seven images (out of 1,000 images) with labels of the Train class. In five of these, the samples are extremely challenging (Figure 7), while in the other two, the labels are just wrong (Figure 8). Analyzing the BDD100K training set, we found only 27 images with labels of the Train class (out of 7,000 images), where the labels were wrong in 13 of them. All this explains the poor results seen in Table V for the Train class. As can be seen in Figures 9 and 10, Cityscapes also have some errors and inconsistencies regarding human labeling. In Figure 9, we see how the co-training method properly labels a traffic light and a traffic sign, while humans did not label them. Since this image is from the validation set of Cityscapes, the self-labels are considered wrong. In the same image, a rider on a bicycle is labeled as a fully visible Person by humans, while the upper part of the person is labeled as a Rider by co-training, and the rest is labeled as a Bicycle. Again, we think that the self-labels better represent this image patch’s semantics. In Figure 10, we can see how different criteria have been used to label Fence instances in different images. 11 shows another example of inconsistency, in this case from a Mapillary Vistas image. We can see how the closed Traffic Light instance has been fully labeled as such (lightbox and its support), while in the further away ones, only the light boxes are labeled as Traffic Light.